8. When would you use the point-slope form to find the equation of a line? Be specific.
Answer #1

We can use point -slope form in two ways

1) if a point through which a line passes is given in question and slope of line that is with what angle the line makes with positive x axis is given then we can find the slope as slope = tan(\Theta) . If they are given than it can be used directly using formula, (y-y1)=m(X-X1), where m= slope and point P=(x1,y1).

2) if two points through which a line passes is given im question that is P1=(x1,y1) and P2 =(x2,y2)  then we can find slope using m=( y2-y1)/(x2-x1). Then we can use the above formula to find equation of line .